A insulating rod is bent into a semicircular arc of radius a, and a total charge Q is distributed uniformly along the rod. Calculate the electric potential at the center of curvature of the arc if the potential is assumed to be zero at infinity.
1
Expert's answer
2021-07-05T17:53:34-0400

Radius=a

Charge=Q

"V=\smallint \frac{dq}{4 \pi\epsilon_0 r}"

"dq=\lambda dl"

"dl=ad\theta"

"\lambda=\frac{\theta}{\pi a}"

"dq=\frac{\theta}{\pi a} (ad\theta)"

"dq=\frac{\theta d\theta}{\pi}"



Limit

"\theta=0°" to "\theta=\pi"

"V=\frac{1}{4\pi\epsilon_0}\int\frac{dq}{a}"

"V=\frac{1}{4\pi\epsilon_0}\smallint _{0}^ \pi \frac{Q d\theta}{\pi a}"

"V=\frac{Q}{4 \pi\epsilon_0 a}"